# DINO-Foresight: Looking into the Future with DINO

### v1

@@ Abstract
Predicting future dynamics is crucial for applications like autonomous driving and robotics, where understanding the environment is key. Existing pixel-level methods are computationally expensive and often focus on irrelevant details. To address these challenges, we introduce DINO-Foresight, a novel framework that operates in the semantic feature space of pretrained Vision Foundation Models (VFMs). Our approach trains a masked feature transformer in a self-supervised manner to predict the evolution of VFM features over time. By forecasting these features, we can apply off-the-shelf, task-specific heads for various scene understanding tasks. In this framework, VFM features are treated as a latent space, to which different heads attach to perform specific tasks for future-frame analysis. Extensive experiments show that our framework outperforms existing methods, demonstrating its robustness and scalability. Additionally, we highlight how intermediate transformer representations in DINO-Foresight improve downstream task performance, offering a promising path for the self-supervised enhancement of VFM features. We provide the implementation code at https://github.com/Sta8is/DINO-Foresight.
@@ 1 Introduction
Predicting future states in video sequences is a key challenge in computer vision and machine learning, with important applications in autonomous systems such as self-driving cars and robotics [26, 22], which must navigate dynamic environments safely. While advances have been made in video understanding [44, 56] and generation [38], predicting future states remains difficult, especially in complex scenarios involving interactions between multiple objects over long time horizons.
Most existing approaches focus on future prediction at the pixel level, aiming to model both low-level appearance and semantic changes. These methods include generative approaches like diffusion-based models [99, 27, 38, 39, 34, 8], autoregressive models [42, 40, 52, 86], and masked video generation [95, 96]. However, for decision-making systems like self-driving cars, the primary goal is not to generate realistic-looking RGB future frames but to perceive a scene and predict its evolution over time at a more semantic or abstract level, such as identifying what objects exist in the scene and where. Thus, future prediction approaches defined at the pixel space may allocate excessive capacity to modeling low-level appearance, which may be irrelevant to the scene understanding needed for decision-making.
Previous work has shown that directly predicting task-specific features for future frames can be more effective than generating full RGB frames and then performing tasks like segmentation or depth estimation [62]. However, existing methods have limitations: some are designed for a single task (e.g., segmentation) [66, 18, 79], making it difficult to handle multiple tasks simultaneously, while others attempt to predict features for multiple tasks at once [41], resulting in complex architectures that do not scale well.
To address these challenges, we leverage recent advances demonstrating the effectiveness of features derived from large-scale Vision Foundation Models (VFMs) for diverse scene understanding tasks, with strong generalization across varied [9, 68, 69]. Building on this foundation, in this work we propose a novel approach to future prediction that operates directly within the space of VFM features. Specifically, we introduce a masked feature transformer architecture, which processes VFM features from previous frames to predict their future counterparts. Using these features, our method inherits a strong and versatile understanding of the scene, offering a unified semantic representation of future states. This representation can be readily adapted to various tasks through task-specific prediction heads applied to the predicted features (see Fig. 1).
Our method offers several key advantages:
Simplicity, scalability, and semantic focus:
It leverages the benefits of generative pixel-level methods, providing simplicity, self-supervised training, and scalability, while avoiding the need to model low-level appearance details.
Task-agnostic:
Unlike prior task-specific semantic future prediction methods, our approach is flexible and can be applied to any scene understanding task, without retraining the model when task requirements change.
Plug-and-play modular task heads:
We use a modular approach where different task-specific heads can be easily added or removed from the framework without retraining the future feature prediction model.
Our contributions are fourfold: (1) We introduce DINO-Foresight, a self-supervised method for semantic future prediction that leverages VFM features. This approach is scalable and can be applied to a variety of scene understanding tasks in a straightforward and practical way. (2) We develop an encoder-decoder transformer architecture for VFM feature forecasting, using a pretrained VFM as the encoder and a masked feature transformer as the decoder that performs the next-frame feature prediction task (see Fig. 1). This setup efficiently propagates multi-layer and high-resolution features, which is crucial for scene understanding task performance. (3) We demonstrate the effectiveness of our approach through extensive experiments, showing superior performance in multiple scene understanding tasks, including semantic and instance segmentation, depth estimation, and surface normals prediction. (4) Last, but not least, we show that the intermediate features within our masked transformer can further improve the performance of downstream tasks. This finding illustrates the promise of our method as a self-supervised visual learning strategy that enhances the already strong VFM features.
@@ 2 Related Work
Figure 1: Future Prediction with Masked Feature Transformer. Our framework trains a masked transformer to predict the temporal evolution of features extracted by a frozen visual encoder based on context frames in a self-supervised manner, minimizing SmoothL1 loss between predicted and future features. By forecasting these rich and versatile features, task-specific prediction heads—such as semantic segmentation, depth, and surface normals—can be effortlessly employed at test time, enabling modular and efficient scene understanding.
@@
Future Prediction  Video future prediction and generation hold significant promise for applications like autonomous driving and robotics, where crafting future frames from past observations is crucial yet challenging due to the high dimensionality and unpredictability of video data. Traditional methods like Convolutional Long Short-Term Memory networks (Conv-LSTMs) [66, 91, 87, 10, 54, 89, 28] have grappled with maintaining visual realism and consistency over time. Subsequent advancements have leveraged generative adversarial networks (GANs) and variational autoencoders (VAEs) [92, 84, 10, 53, 3], alongside diffusion models [38, 39, 27, 34], to enhance spatial-temporal coherence and improve the quality of predictions. Furthermore, transformer-based models have also been adapted to videos, utilizing auto-regressive and masked modeling objectives to capture video dynamics [96, 95, 33, 86].
@@
Future Feature Prediction  An emerging approach for future-frame prediction focuses on forecasting intermediate features from an encoder rather than predicting raw RGB values [66, 74, 77, 15, 45, 63, 73, 43, 83, 100]. This strategy models the abstract representations learned by the encoder, which are then used by task-specific heads for downstream tasks. While effective, these methods often rely on encoders specialized for specific tasks or datasets, limiting their generalizability across domains. To address this, we use encoders from visual foundation models, which, with their large-scale pre-training, perform well across a variety of tasks and generalize effectively to new scenes.
Vision Foundation Models (VFMs)
@@
VFMs have transformed computer vision, achieving strong performance across a range of visual tasks. Trained on large-scale datasets, these models learn rich, transferable visual representations. Notable examples include DINOv2 [9, 68], a self-supervised model based on self-distillation; CLIP and its variants [69, 24, 78, 25], which align visual representations with natural language; and SAM (Segment Anything Model) [50], a foundation model for object segmentation. In this work, we explore the potential of VFM features for semantic future prediction tasks, connecting static visual understanding with dynamic prediction.
@@
Multi Task Learning  Multi-Task Learning (MTL) is a learning paradigm that enables simultaneous training of models on multiple related tasks [64, 65, 67, 82], promoting shared representations and improving performance across tasks. Traditional MTL frameworks often use parameter sharing [48, 75, 5] or task interaction allowing exchange of information [7, 14, 16, 65, 71]. Other approaches employ a strategy that incrementally increases the model’s depth during training, enabling the network to learn task-specific representations in a more resource-efficient way [1, 19, 32, 61, 97]. Recently, the emergence of large-scale pretrained models has led to the introduction of adapter-based multi-task fine-tuning approaches [55, 59]. In our work, we leverage VFM features to provide a unified, scalable and modular framework for future prediction. Our approach enables seamless integration of multiple tasks without retraining or complex adaptations.
@@ 3 Methodology
Our semantic future prediction framework leverages VFMs trained on large-scale data, which have proven highly effective in various scene understanding tasks and in generalizing to unseen scenes. We propose training a masked feature transformer model in a self-supervised manner to predict the evolution of these VFM features across future frames. By forecasting these features, we can apply various off-the-shelf, task-specific heads for scene understanding. In this approach, we treat the feature representations of VFMs as a latent space, to which different heads can be connected to perform specific tasks for future-frame analysis.
Fig. 1 provides an overview of our approach, with the key components detailed in the following sections: Section 3.1 describes how target features are generated from multi-layer VFM features and the strategies used to manage their high dimensionality. Section 3.2 explains how we formulate the future-frame prediction task as a masked feature modeling problem and the transformer architecture used. Section 3.3 covers techniques for efficiently training the transformer to predict high-resolution VFM features. Finally, Section 3.4 presents our modular framework for multi-task predictions using future-frame features and details how prediction heads are trained and integrated.
@@ 3.1 Hierarchical Target Feature Construction
In Fig. 2, we provide an overview of how the target feature space for the masked feature transformer is constructed. Below, we outline the main steps involved.
Multi-Layer VFM Feature Extraction  Scene understanding models often benefit from processing features across multiple layers of an image encoder [70, 60, 57, 17, 98, 12], especially when the encoder is frozen, as in our approach. To fully leverage the pretrained representations of VFMs, we propagate features extracted from multiple layers of the VFM.
Our framework uses VFMs based on the Vision Transformer (ViT) architecture, though the approach can be extended to other architectures. Given a sequence of $N$ image frames $\mathbf{X}\in\mathbb{R}^{N\times H^{\prime}\times W^{\prime}\times 3}$, let $\mathbf{F}^{(l)}\in\mathbb{R}^{N\times H\times W\times D_{enc}}$ represent the features extracted from layer $l$ of the ViT model. Here, $D_{enc}$ is the feature dimension, and $H\times W$ is the spatial resolution, which are consistent across layers in ViT-based models. To form the target feature space on which the masked feature transformer operates, we concatenate the features from $L$ layers along the channel dimension, resulting in $\mathbf{F}_{\text{concat}}\in\mathbb{R}^{N\times H\times W\times L\cdot D_{enc}}$. These concatenated features capture rich semantic information from the input images at multiple levels of abstraction.
Dimensionality Reduction
The concatenated features $\mathbf{F}_{\text{concat}}$ have high dimensionality, so we apply dimensionality reduction to simplify the prediction task for the transformer while retaining essential information. In this work, we use Principal Component Analysis (PCA) to reduce the dimensionality, transforming $\mathbf{F}_{\text{concat}}$ into a lower-dimensional representation $\mathbf{F}_{\text{PCA}}\in\mathbb{R}^{N\times H\times W\times D}$, where $D\ll L\cdot D_{enc}$. These PCA-reduced features, $\mathbf{F}_{\text{PCA}}$, serve as the target features on which the masked feature transformer operates, i.e., $\mathbf{F}_{\text{TRG}}=\mathbf{F}_{\text{PCA}}$. While PCA is used in this work, alternative methods, such as Variational Autoencoders (VAE), could also be explored in future work.
@@ 3.2 Self-Supervised Future Feature Prediction with Masked Transformers
Figure 2: Hierarchical Target Feature Construction for Masked Feature Transformer. Our framework builds a feature space by extracting and concatenating multi-layer features from a frozen ViT encoder, capturing semantic information at different levels of abstraction. To reduce their high dimensionality, we apply PCA, creating more compact features for the masked feature transformer to operate on.
Masked Feature Transformer Architecture
@@
Inspired by previous video generation models [11, 33], we implement the future feature prediction task using a self-supervised masked transformer architecture. The task involves predicting future frames in a video sequence consisting of $N$ frames, where $N_{c}$ are context frames and $N_{p}$ are future frames to be predicted, such that $N=N_{c}+N_{p}$. Given the target features $\mathbf{F}_{\text{TRG}}\in\mathbb{R}^{N\times H\times W\times D}$ for these $N$ frames, the future-frame tokens are masked, and the transformer must predict these missing tokens by processing all the tokens from the entire sequence, i.e., all the $N\cdot H\cdot W$ tokens.
@@
The transformer architecture begins with a token embedding stage, where each token is projected from $D$ dimensions into the transformer’s hidden dimension $D_{dec}$ through a linear layer. During training, the tokens corresponding to the future frames are replaced with a learnable $D_{\text{dec}}$-dimensional [MASK] vector. During inference, these [MASK] tokens are appended after the context frames. Each token also receives a position embedding to retain both temporal and spatial information across the sequence.
@@
The tokens are then passed through a series of transformer layers. Standard self-attention layers in transformers have quadratic time complexity with respect to the number of tokens, making them computationally expensive for high-resolution, multi-frame sequences. To address this, we follow the approach from recent video transformers [2, 33] and decompose the attention mechanism into temporal and spatial components. Temporal attention is applied across tokens with the same spatial position in different frames, capturing the dynamics over time. Spatial attention, on the other hand, operates within individual frames, focusing on spatial interactions. Thus, each transformer layer consists of a temporal Multi-Head Self-Attention (MSA) layer, a spatial MSA layer, and a feedforward MLP layer. After passing through the transformer layers, a linear prediction layer maps the output token embeddings from the hidden dimension $D_{\text{dec}}$ back to the feature dimension $D$, producing the predicted feature map $\mathbf{\tilde{F}}_{\text{TRG}}\in\mathbb{R}^{N\times H\times W\times D}$.
Training Objective for Masked Feature Modeling
@@
We frame the future-frame prediction as a continuous regression problem and optimize a self-supervised training objective based on the SmoothL1 loss between the predicted features $\mathbf{\tilde{F}}_{\text{TRG}}$ and the ground truth features $\mathbf{F}_{\text{TRG}}$ at the masked locations.
The loss function is defined as:
@@
$\mathcal{L}_{\mathrm{MFM}}=\underset{x\in\mathcal{X}}{\mathbb{E}}\left[\sum_{p\in\mathcal{P}}M(p)\cdot L\left(\mathbf{F}_{\text{TRG}}(p),\mathbf{\tilde{F}}_{\text{TRG}}(p)\right)\right],$ (1)
@@
where $\mathcal{X}$ denotes the training dataset, $\mathcal{P}=[1:N]\times[1:H]\times[1:W]$ represents the set of positions across the $N$, $H$, and $W$ dimensions, $\mathbf{F}_{\text{TRG}}(i)$ and $\mathbf{\tilde{F}}_{\text{TRG}}(i)$ are the ground truth and predicted feature vectors at position $p$, respectively, and $M$ is a binary mask that is non-zero only at the positions corresponding to future frames. The function $L(\cdot,\cdot)$ computes the SmoothL1 loss between two feature vectors:

### v2

@@ Abstract
Predicting future dynamics is crucial for applications like autonomous driving and robotics, where understanding the environment is key. Existing pixel-level methods are computationally expensive and often focus on irrelevant details. To address these challenges, we introduce DINO-Foresight, a novel framework that operates in the semantic feature space of pretrained Vision Foundation Models (VFMs). Our approach trains a masked feature transformer in a self-supervised manner to predict the evolution of VFM features over time. By forecasting these features, we can apply off-the-shelf, task-specific heads for various scene understanding tasks. In this framework, VFM features are treated as a latent space, to which different heads attach to perform specific tasks for future-frame analysis. Extensive experiments show the very strong performance, robustness and scalability of our framework. Project page and code at https://dino-foresight.github.io/.
@@ 1 Introduction
Predicting future states in video sequences is a key challenge in computer vision and machine learning, with important applications in autonomous systems like self-driving cars and robotics (Finn et al., 2016; Dosovitskiy and Koltun, 2017). These systems must navigate dynamic environments safely, yet predicting future states remains difficult—especially in complex scenarios involving multi-object interactions over long time horizons.
Recent approaches focus on generating realistic RGB future frames using latent generative modeling (Rombach et al., 2022). These methods first compress RGB data into a latent space, such as continuous (Rombach et al., 2022) or discrete (Esser et al., 2021) Variational Autoencoder (VAE) representations. Then, they train generative models—like diffusion models (Zheng et al., 2024; Gao et al., 2024; Ho et al., 2022a, b; Harvey et al., 2022; Brooks et al., 2024), autoregressive models (Hu et al., 2023; Hong et al., 2023; Kondratyuk et al., 2024; Wang et al., 2024), or masked video generation (Yu et al., 2023, 2024)—to predict future states in this compressed space. While this reduces dimensionality and improves training stability (Rombach et al., 2022), VAE latents often lack semantic alignment, making them hard to interpret or use in downstream scene understanding tasks (see Table 2). Moreover, these methods must model both low-level appearance and high-level semantics, even though decision-making systems (e.g., self-driving cars) primarily need semantic scene understanding—what objects exist and where they are. Latent generative approaches may waste capacity on irrelevant low-level details, compromising temporal semantic accuracy.
In contrast, vision foundation models (VFMs) have revolutionized scene understanding with robust, transferable features (Oquab et al., 2024; Radford et al., 2021; Venkataramanan et al., 2025; Bardes et al., 2024). This raises a key question: can VFM features serve as a semantically rich, high-dimensional latent space for precise future prediction?
In this work, we propose precisely this idea. Instead of predicting future low-level VAE latents, we forecast the temporal evolution of VFM features directly. This shift brings several important advantages: (a) Beyond low-level latents. It leverages semantically meaningful representations, inheriting strong scene understanding. (b) Dynamic semantics over raw frames. It avoids full-frame synthesis, letting the model focus on meaningful dynamics, reducing complexity and sidestepping challenges like multimodal pixel distributions. (c) Scalable multi-task support and modular integration. Forecasted features align with downstream tasks, enabling plug-and-play integration with pretrained classifiers, segmenters or task-specific heads—without the need to retrain the core feature predictor (see Figure 1). This represents a significant departure from prior semantic feature prediction methods (Luc et al., 2017), which face significant challenges with multi-task scalability. Such approaches either require training a separate model for each task (Nabavi et al., 2018; Chiu et al., 2020; Terwilliger et al., 2019) or involve predicting features from multiple task-specific models simultaneously (Hu et al., 2020; Karypidis et al., 2025), resulting in more complex and less scalable architectures.
Figure 1: Forecasting VFM Features for Future Frames. At the core of our approach is the prediction of future VFM feature evolution. To this end, we train a masked transformer model in a self-supervised manner to forecast these features from context frames, minimizing SmoothL1 loss between predicted and actual future features. By forecasting these rich and versatile features, task-specific prediction heads—such as semantic segmentation, depth, and surface normals—can be effortlessly employed at test time, enabling modular and efficient multi-task scene understanding.
In this work we show that forecasting high-dimensional VFM features is not only feasible but also achieves strong performance, offering a new path toward semantically grounded future prediction.
In summary, our contributions are: (1) We introduce DINO-Foresight, a self-supervised method for semantic future prediction that builds on the key idea of forecasting VFM features—our core contribution. Unlike latent generative methods that rely on low-level VAE features, our approach delivers precise future scene understanding without modeling unnecessary appearance details, enabling a unified model for diverse scene understanding tasks. (2) To realize this idea, we design an efficient masked feature transformer (see Figure 1) that propagates multi-layer, high-resolution features critical for achieving strong task performance. (3) Experimental results demonstrate a unique advantage of our approach - our single model successfully handles multiple future-frame understanding tasks (semantic segmentation, instance segmentation, depth prediction, and surface normal prediction) where previous approaches required multiple specialized models. Furthermore, as we show in Appendix Subsection A.2 intermediate features within our masked transformer can further improve downstream task performance, highlighting its potential as a self-supervised learning strategy that enhances the already strong VFM features.
@@ 2 Related Work
Video Prediction  is an extensively studied problem. Early approaches based Convolutional LSTMs (Nabavi et al., 2018; Xu et al., 2018; Wang et al., 2018; Castrejon et al., 2019; Lee et al., 2021; Wu et al., 2021; Gao et al., 2022) struggled to maintain both visual quality and temporal consistency. Subsequent developments introduced generative adversarial networks and variational autoencoders (Yan et al., 2021; Vondrick et al., 2016b; Castrejon et al., 2019; Lee et al., 2018; Babaeizadeh et al., 2018), alongside diffusion models (Ho et al., 2022a, b; Gao et al., 2024; Harvey et al., 2022), to enhance spatial-temporal coherence and improve the quality of predictions. Furthermore, transformer-based models have also been adapted to videos, utilizing auto-regressive and masked modeling objectives to capture video dynamics (Yu et al., 2024, 2023; Gupta et al., 2023; Wang et al., 2024).
Latent Generative Models  Current state-of-the-art video prediction methods (Gupta et al., 2023; Yu et al., 2023, 2024; Gao et al., 2024) build upon latent generative approaches (Rombach et al., 2022). These frameworks employ an autoencoder to compress RGB frames into a latent space, then train generative models to predict future states in this compressed representation. While sharing superficial similarities with our approach—including the use of latent spaces and masked transformers in some cases (Gupta et al., 2023; Yu et al., 2023)—these methods differ fundamentally. Their latent spaces primarily encode low-level visual information, requiring reconstruction back to RGB space and forcing the model to simultaneously handle both appearance details and semantic changes.
Our key innovation lies in forecasting VFM features that natively encode high-level semantic information, enabling direct application of task-specific prediction heads. As demonstrated in Table 2, conventional VAE latent spaces cannot match this capability. Our experimental results in Table 1 further show these methods’ limitations in predicting semantic scene evolution, highlighting the distinct advantages of VFM feature forecasting for scene understanding tasks. Concurrently to our work, DINO-WM (Zhou et al., 2025), also leverages DINOv2 features for world modeling with action-conditioned planning in simulated environments, while our work targets multi-task dense semantic forecasting in real world scenarios.
@@
Semantic Future Prediction  An emerging approach for future-frame prediction focuses on forecasting intermediate features from an encoder rather than raw RGB values (Nabavi et al., 2018; Lin et al., 2021; Saric et al., 2020; Sun et al., 2019; Chen and Han, 2019; Jin et al., 2017; Luc et al., 2018; Šarić et al., 2019; Hu et al., 2021; Vondrick et al., 2016a; Zhong et al., 2023). This strategy models abstract encoder representations, which task-specific heads use for downstream tasks. Early methods in this paradigm include F2F (Luc et al., 2018), which regresses Mask-RCNN’s feature pyramid, and F2MF (Saric et al., 2020), which improves feature prediction using flow-based warping. APANet (Hu et al., 2021) aggregates multi-level features via an auto-path mechanism for instance segmentation, while PFA (Lin et al., 2021) enhances global structures and suppresses local details for more predictable features. Recently, Futurist (Karypidis et al., 2025) introduced a multi-modal semantic forecasting approach for semantic segmentation and depth maps. While effective, these methods often rely on task- or dataset-specific encoders, limiting practicality and scalability. To address this, we use VFM encoders, which, due to large-scale pre-training, perform well across diverse tasks and generalize effectively to new scenes without retraining.
@@
Vision Foundation Models (VFMs)  VFMs have transformed computer vision, achieving strong performance across a range of visual tasks. Trained on large-scale datasets, these models learn rich, transferable visual representations. Notable examples include DINOv2 (Caron et al., 2021; Oquab et al., 2024), a self-supervised model based on self-distillation; Franca (Venkataramanan et al., 2025), a fully-open VFM for scalable self-supervised representation learning; CLIP and its variants (Radford et al., 2021; Fang et al., 2023; Sun et al., 2023; Fang et al., 2024), which align visual representations with natural language; SAM (Segment Anything Model) (Kirillov et al., 2023), a foundation model for image segmentation and V-JEPA(Bardes et al., 2024), a video representation learning method via feature prediction in latent space. In this work, we explore the potential of VFM features for semantic future prediction tasks, connecting static visual understanding with dynamic prediction.
@@
Multi-Task Learning  Multi-Task Learning (MTL) is a learning paradigm that enables simultaneous training of models on multiple related tasks (Maninis et al., 2019; Misra et al., 2016; Neseem et al., 2023; Vandenhende et al., 2022), promoting shared representations and improving performance across tasks. Traditional MTL frameworks often use parameter sharing (Kendall et al., 2018; Sener and Koltun, 2018; Bekoulis et al., 2018) or task interaction allowing exchange of information (Bragman et al., 2019; Chen et al., 2023a, b; Misra et al., 2016; Ruder et al., 2019). Other approaches employ a strategy that incrementally increases the model’s depth during training, enabling the network to learn task-specific representations in a more resource-efficient way Aich et al. (2023); Choi and Im (2023); Guo et al. (2020); Lu et al. (2017); Zhang et al. (2022). Recently, the emergence of large-scale pretrained models has led to the introduction of adapter-based multi-task fine-tuning approaches (Liang et al., 2022; Liu et al., 2022). In our work, we leverage VFM features to provide a unified, scalable and modular framework for future prediction. Our approach enables seamless integration of multiple tasks without retraining or complex adaptations.
@@ 3 Methodology
Our semantic future prediction approach builds on forecasting VFM features – powerful representations that excel at scene understanding tasks and generalize well to unseen environments. We realize this through a masked transformer model trained via self-supervision to predict the temporal evolution of VFM features. This forecasted feature space serves as a latent representation that can flexibly connect to various off-the-shelf task-specific heads for future-frame analysis.
Figure 1 provides an overview of our approach, with the key components detailed in the following sections: Section 3.1 describes how target features are generated from multi-layer VFM features. Section 3.2 presents our formulation of VFM feature forecasting as a masked feature modeling problem and details the model architecture. Section 3.3 covers efficient training techniques for high-resolution VFM feature prediction. Finally, Section 3.4 introduces our modular framework for multi-task future-frame analysis and details how prediction heads are trained and integrated.
@@ 3.1 Hierarchical Target Feature Construction
In Figure 2, we provide an overview of how the target feature space for the feature prediction model is constructed. Below, we outline the main steps involved.
Figure 2: Hierarchical Target Feature Construction for the Feature Prediction Model. Our framework constructs a feature space by extracting and concatenating multi-layer features from a frozen ViT encoder, capturing semantic information at varying abstraction levels. PCA is applied to reduce dimensionality, creating compact features.
Multi-Layer VFM Feature Extraction  Scene understanding models often benefit from processing features across multiple layers of an image encoder (Ranftl et al., 2021; Long et al., 2015; Lin et al., 2017; Cheng et al., 2022; Zhao et al., 2017; Chen et al., 2017), especially when the encoder is frozen, as in our approach. To fully leverage the pretrained VFMs, we propagate features extracted from multiple layers of the VFM. We ablate the impact of multi-layer features in Appendix Subsection A.4.
Our framework uses VFMs based on the Vision Transformer (ViT) architecture, though the approach can be extended to other architectures. Given a sequence of $N$ image frames $\mathbf{X}\in\mathbb{R}^{N\times H^{\prime}\times W^{\prime}\times 3}$, let $\mathbf{F}^{(l)}\in\mathbb{R}^{N\times H\times W\times D_{enc}}$ represent the features extracted from layer $l$ of the ViT model. Here, $D_{enc}$ is the feature dimension, and $H\times W$ is the spatial resolution, which are consistent across layers in ViT-based models. To form the target feature space on which the feature prediction model operates, we concatenate the features from $L$ layers along the channel dimension, resulting in $\mathbf{F}_{\text{concat}}\in\mathbb{R}^{N\times H\times W\times L\cdot D_{enc}}$. These concatenated features capture rich semantic information from the input images at multiple levels of abstraction.
Dimensionality Reduction  The concatenated features $\mathbf{F}_{\text{concat}}$ have high dimensionality, so we apply dimensionality reduction to simplify the feature prediction task while retaining essential information. In this work, we use Principal Component Analysis (PCA) to reduce the dimensionality, transforming $\mathbf{F}_{\text{concat}}$ into a lower-dimensional representation $\mathbf{F}_{\text{PCA}}\in\mathbb{R}^{N\times H\times W\times D}$, where $D\ll L\cdot D_{enc}$. These PCA-reduced features, $\mathbf{F}_{\text{PCA}}$, serve as the target features on which the future feature prediction operates, i.e., $\mathbf{F}_{\text{TRG}}=\mathbf{F}_{\text{PCA}}$. In Appendix Subsection A.1, we show that this PCA-based compression retains essential information and enhances downstream performance.
@@ 3.2 Self-Supervised Future Feature Prediction with Masked Transformers
Masked Feature Transformer Architecture  Inspired by previous video generation models (Chang et al., 2022; Gupta et al., 2023; Karypidis et al., 2025), we implement the future feature prediction task using a self-supervised masked transformer architecture. The task involves predicting future frames in a video sequence consisting of $N$ frames, where $N_{c}$ are context frames and $N_{p}$ are future frames to be predicted, such that $N=N_{c}+N_{p}$. Given the target features $\mathbf{F}_{\text{TRG}}\in\mathbb{R}^{N\times H\times W\times D}$ for these $N$ frames, the future-frame tokens are masked, and the transformer must predict these missing tokens by processing all the tokens from the entire sequence, i.e., all the $N\cdot H\cdot W$ tokens.
@@
The architecture begins with a token embedding stage, where each token is projected from $D$ dimensions into the transformer’s hidden dimension $D_{dec}$ through a linear layer. During training, the tokens corresponding to the future frames are replaced with a learnable $D_{\text{dec}}$-dimensional [MASK] vector. During inference, these [MASK] tokens are appended after the context frames. Each token also receives a position embedding to retain both temporal and spatial information across the sequence.
@@
The tokens are then passed through a series of transformer layers. Standard self-attention layers in transformers have quadratic time complexity with respect to the number of tokens, making them computationally expensive for high-resolution, multi-frame sequences. To address this, we follow the approach from recent video transformers (Arnab et al., 2021; Gupta et al., 2023; Karypidis et al., 2025) and decompose the attention mechanism into temporal and spatial components. Temporal attention is applied across tokens with the same spatial position in different frames, capturing the dynamics over time. Spatial attention, on the other hand, operates within individual frames, focusing on spatial interactions. Thus, each transformer layer consists of a temporal Multi-Head Self-Attention (MSA) layer, a spatial MSA layer, and a feedforward MLP layer. After passing through the transformer layers, a linear prediction layer maps the output token embeddings from the hidden dimension $D_{\text{dec}}$ back to the feature dimension $D$, producing the predicted feature map $\mathbf{\tilde{F}}_{\text{TRG}}\in\mathbb{R}^{N\times H\times W\times D}$.
@@
Training Objective  We frame the future-frame prediction as a continuous regression problem and optimize a self-supervised training objective based on the SmoothL1 loss between the predicted features $\mathbf{\tilde{F}}_{\text{TRG}}$ and the ground truth features $\mathbf{F}_{\text{TRG}}$ at the masked locations. The loss is defined as:
@@
$\mathcal{L}_{\mathrm{MFM}}=\underset{x\in\mathcal{X}}{\mathbb{E}}\left[\sum_{p\in\mathcal{P}}L\left(\mathbf{F}_{\text{TRG}}(p),\mathbf{\tilde{F}}_{\text{TRG}}(p)\right)\right],$ (1)
@@
where $\mathcal{X}$ denotes the training dataset, $\mathcal{P}=[N_{c}+1:N]\times[1:H]\times[1:W]$ represents the set of positions across the $H\times W$ spatial dimensions of the $N_{p}$ future frames, and $\mathbf{F}_{\text{TRG}}(p)$ and $\mathbf{\tilde{F}}_{\text{TRG}}(p)$ are the ground truth and predicted feature vectors at position $p$, respectively. $L(\cdot,\cdot)$ computes the SmoothL1 loss between two feature vectors: